# Haptic bilateral teleoperation system for free-hand dental procedures

### v2

@@ V Discussions and conclusions
This paper proposed a FH-HBTS for free-hand dental procedures. The system features visual and haptic feedback, that are appropriately combined and coordinated to allow for a natural teleoperation experience. To improve the visibility of the intervention site, we designed a dedicated end-effector accommodating an eye-in-hand endoscopic camera, and ensuring compatibility with existing clinical tools. We also improved the accuracy of the procedure by introducing human tremor filtering and dynamic scaling of velocities. By providing the eye-hand coordination and the TCP control, we enabled the human operator to perceive the motion correspondence between their movements and those of the end-effector, observed in the camera view, preserving manual dexterity and expertise. Additionally, by appropriately transforming the virtualized force feedback, we also ensured the force correspondence, enabling the human operator to experience a highly natural sensory perception. Finally, we enabled safe interactions, by scaling the motion references provided to the admittance controller. The proposed strategy allows for accurate force limitation in any contact state and for any orientation of the interaction tool, without requiring any prior modeling of the intervention site.
@@
We believe that our FH-HBTS represents a crucial step towards a broader adoption of robotics in dentistry, as it allows for accurate treatments, while preserving the operator’s cognitive and technical abilities. With respect to manual procedures, our system allows to preserve operator’s ergonomics, track operator’s performance, and enable remote healthcare.
@@
We emphasize that the proposed system allows to collect haptic and motion data. While this feature is crucial during the training phase of new operators for the evaluation of learning metrics, i.e learning curves [56], it is also central towards the deployment of data-driven control architectures that leverage demonstrations to synthesize control policies [57]. Finally, these data can also be used to learn cost functions [58], and hence intentions, of experts fulfilling a given task. In the context of dental procedures, this means that data collected from experts executing the procedure can be used to train policies for autonomous robots.
With our future work we plan to extend the FH-HBTS with autonomous patient motion compensation to further enhance the system’s safety, and to benchmark the performance against the da Vinci Surgical SystemTM. Also, we plan to conduct experiments with professionals and active dental tools both to confirm the system usability under more representative clinical conditions and to collect expert policies for (semi-)autonomous procedures.
@@ Acknowledgments
The authors would like to thank Sabatino Panico for contributing to the system calibration and experimental campaigns.

### v3

@@ VI Conclusions and Future Perspectives
This paper proposed a FH-HBTS for free-hand dental procedures. The system features visual and haptic feedback, that are appropriately combined and coordinated to allow for a natural teleoperation experience. To improve the visibility of the intervention site, we designed a dedicated end-effector accommodating an eye-in-hand endoscopic camera, and ensuring compatibility with existing clinical tools. We also improved the accuracy of the procedure by introducing human tremor filtering and dynamic scaling of velocities. Through eye-hand coordination and TCP control, we enabled the operator to perceive the motion correspondence between their movements and those of the end-effector, observed in the camera view, preserving manual dexterity and expertise. Additionally, by appropriately transforming the virtualized force feedback, we also ensured the force correspondence, enabling the human operator to experience a highly natural sensory perception. Finally, we enabled safe interactions by scaling the motion references provided to the admittance controller. The proposed strategy allows for accurate force limitation in any contact state and for any orientation of the interaction tool, without requiring any prior modeling of the intervention site.
@@
We believe that our FH-HBTS represents a crucial step towards a broader adoption of robotics in dentistry, as it allows for accurate treatments, while preserving the operator’s cognitive and technical abilities. With respect to manual procedures, our system allows improving operator’s ergonomics, tracking operator’s performance, and perspectively enabling remote healthcare.
@@
We emphasize that the proposed system allows collecting haptic and motion data. While this feature is crucial during the training phase of new operators for the evaluation of learning metrics, i.e learning curves [71], it is also central towards the deployment of data-driven control architectures that leverage demonstrations to synthesize autonomous control policies [19]. Finally, these data can also be used to learn cost functions [21], and hence intentions, of experts fulfilling a given task. In the context of dental procedures, this means that data collected from experts executing the procedure can be used to train policies for autonomous robots.
With our future work we plan to extend the FH-HBTS with autonomous patient motion compensation to further enhance system’s safety and felxibility, and to benchmark the performance against the da Vinci Surgical SystemTM. In addition, future experimental testbeds will be designed to progressively incorporate clinically relevant sources of complexity that are not yet addressed in the present study, including constrained inter-proximal access, interaction with soft tissues, fluid presence, and limited workspace conditions. In this context, we plan to collect expert demonstrations for the development of (semi-)autonomous dental procedures. Beyond algorithmic extensions, a key direction of future work concerns the evolution of the experimental platform itself, which will be redesigned towards a fully engineered, application-driven system, explicitly accounting for the mechanical, sensing, and ergonomic requirements of dental procedures, and addressing the hardware-related limitations identified in the present study.
@@ Acknowledgments
The authors would like to thank X for their assistance during the in-vitro pre-clinical validation and Y for supporting the EHCC implementation and end-to-end system validation, making this work possible.